OpenStudy (anonymous):

Solve tan2x = sqrt3/3

OpenStudy (anonymous):

tan(2x)=sqrt3 or, 2tanx/(1-tan^2(x)) = sqrt(3) or, 2tanx = sqrt(3) -sqrt(3) tan^2(x) or, sqrt(3) tan^2(x) +2tan(x) -sqrt(3) =0 or, tanx = -sqrt(3), 1/sqrt(3) if tanx = -sqrt3 , x= 2pi/3 if tanx = 1/sqrt3 x= pi/6 Source: http://answers.yahoo.com/question/index?qid=20100819024801AA2XG0k